OpenStudy (anonymous):

Solve 2/(2x+5)+3/(2x-5)=(10x+5)/(4x^2-25)

OpenStudy (anonymous):

\[\frac{2}{(2x+5)}+\frac{3}{(2x-5)}=\frac{(10x+5)}{(4x^2-25)}\]

OpenStudy (anonymous):

I got x=5/6

OpenStudy (anonymous):

yup , to check it just substitute the x in the equation if the left and right side of the equation are equal therefore you are correct

OpenStudy (anonymous):

oh okay, thanks I forgot about doing that...